# From “The Cliff” to “Virgil”: Mapping the Spectral Diversity of Little Red Dots with JWST/NIRSpec

### v1

@@ 4.5 Color Dependence of the Emission Line FWHMs
The colormap suggests a positive correlation, with the reddest LRDs showing the highest FWHMs. A linear fit yields a statistically significant trend, albeit with substantial scatter (Pearson , ; Spearman , ). The median FWHM, binned by rest-frame color as in Figure 4, declines from 2327 760 km s-1 for [0.30.9 m] 3, to 2067 877 km s-1 for [0.30.9 m] , and 1677 955 km s-1 for [0.30.9 m] 2. This trend agrees with the results of greene23, who found broader lines in redder LRDs (e.g., F277WF444W 1.6) in the UNCOVER field.
Despite this general trend, the scatter is large, and there are notable exceptions. Some blue LRDs exhibit very broad lines—for example, GOODS-N-1001830 from juodzbalis24a and UNCOVER-38018 from greene23 both show FWHM 4000 km s-1. Conversely, some of the reddest LRDs, such as RUBIES-UDS-16053 (taylor25) and “The Cliff” (degraaff25), exhibit narrower lines with FWHM 1500 km s-1.
4.6 LRDs with narrow absorption in the broad emission lines
@@
4.7 Color Dependence of Emission-line Ratios
@@
The physical origin of the large Balmer decrement in LRDs remains poorly understood. Early interpretations invoking heavy obscuration (e.g., kocevski23, greene23) are disfavored in light of the BH⋆ models, which predict more complex radiative transfer effects. In this context, the elevated broad H /H ratios may be boosted by collisional excitation and resonant scattering in the dense, low-metallicity (and low-extinction) gas cocoon surrounding the accreting black hole (e.g., chang25; torralba25b). Supporting this scenario, the LRDs with the most extreme H /H ratios are also those that exhibit narrow self-absorption features superimposed on their broad Balmer lines.
@@
To further test whether narrow emission lines arise from the host galaxy while the optical continuum is AGN-dominated, we examine the correlation between equivalent width (EW) and [0.30.9 m] color for the most prominent narrow lines. The right panel of Figure 7 shows the EW of [O iii]5007 and the of blended H +[O iii]50074364, the latter serving as a proxy for the complex of lines near the Balmer-break ([O ii], [Ne iii], H ) commonly seen in high-EW galaxies. Among these, H +[O iii]50074364 is typically the strongest and most reliably detected, with 45% of LRDs showing S/N 5 in this feature, compared to 20%, 18%, and 1% for [O iii]5007, [Ne iii], and [O ii], respectively. Both EWs show a clear anti-correlation with [0.30.9 m] color: bluer LRDs exhibit stronger narrow-line emission, while redder LRDs show much weaker or undetected lines. This supports the picture that narrow-line features become more prominent toward bluer LRDs, where the host galaxy likely contributes more to the spectrum.
We caution that H +[O iii]50074364 may not be purely narrow, especially in the reddest LRDs where broad H emission can dominate. Still, high-resolution NIRSpec spectra suggest that the narrow H component becomes more prominent in bluer LRDs (e.g., akins24b; juodzbalis24b; rinaldi25), while the broad component dominates over the narrow in redder systems (e.g., deugenio25; jones25; torralba25b) similarly to H .
The UV emission lines in LRDs are generally faint, with few detections of classical AGN tracers such as C iii], C iv, or Mg ii above S/N 5. The exception is Ly, detected in 20% of the sample. While Ly equivalent widths show no clear trend with color, detections are concentrated in bluer LRDs, with a median [0.30.9 m] 2.0. Lowering the detection threshold to S/N 2 boosts the C iii] detection rate to 22%, consistent with it being the brightest UV line after Ly; C iv and Mg ii remain below 5%. Like Ly, C iii] detections span a wide color range but are slightly redder on average, with [0.30.9 m] 2.4—consistent with previous reports of C iii] in some of the reddest LRDs (e.g., labbe24b; deugenio25; torralba25b).
5 A Semi-Empirical Model for LRD SEDs: From “The Cliff” to “Virgil”
5.1 A Two-Component model: galaxy plus BH⋆
To address these constraints, we adopt a semi-empirical model combining a galaxy and a BH⋆ component. Instead of a complex parametric approach using photoionization codes like CLOUDY (cloudy), we use a simplified method where the BH⋆ component is modeled using the observed continuum of the extreme LRD “The Cliff” (degraaff25). A similar strategy was recently adopted by degraaff25b and umeda25, who approximate the BH⋆ SED with a blackbody. Here, we construct the BH⋆ template by fitting a low-order polynomial to the emission-line–free regions of the “The Cliff” spectrum, capturing its shape from the Balmer break to 1 m rest-frame. We assume no UV emission from the BH⋆ component (see § 6 for further discussion). Dust attenuation is applied using an SMC law (pei92), with as the only free parameter modifying the BH⋆ shape. Even modest attenuation can transform the intrinsic curvature of the BH⋆ template into a redder optical continuum that more closely resembles a power law. Varying both the reddening and relative luminosity of the BH⋆ component naturally shifts the SED inflection point to longer wavelengths. Our model does not include an additional IR component to account for potential dust emission from the BH⋆, as in barro25, since the current spectra do not probe the rest-frame NIR (–3 m). Incorporating MIRI data would be essential to assess this contribution and test for warm dust associated with the BH⋆ component.
5.2 Interpreting LRD Diversity Through the Two-Component Model
@@
Our qualitative expectations are confirmed by the distribution of best-fit parameters as a function of rest-frame color across the LRD population. Dividing the sample into three UV-to-optical color bins, [0.30.9 m] , [0.30.9 m] , and [0.30.9 m] , we find that the median stellar attenuation decreases steadily from in the reddest bin to and in the intermediate and bluest bins, respectively. In contrast, the attenuation of the BH⋆ component increases across the same bins, from to and . Meanwhile, the BH⋆ luminosity declines from to and . These trends suggest that redder LRDs are powered by more luminous, less obscured BH⋆ components embedded in more attenuated stellar hosts, while bluer LRDs tend to host fainter, more obscured BH⋆ components and nearly dust-free galaxies.
@@
Overall, the two-component model successfully fits LRDs across the full range of [0.30.9 m] colors and SED types, suggesting that the observed diversity arises primarily from variations in the galaxy-to-BH⋆ luminosity ratio, coupled with modest obscuration of the BH⋆ component. As this ratio decreases toward bluer colors, the crossover between the stellar (blue) and BH⋆ (red) continua shifts to longer wavelengths, from 0.4 m to 0.7 m, delaying the onset of the steep optical rise.
@@
To quantify the shift in the inflection, Figure 10 shows the wavelength at which the BH⋆ and stellar components contribute equally to the total SED. While this crossover does not mark the exact inflection point, the steep slope of the BH⋆ component ensures that the turnover typically occurs slightly blueward of this intersection. The distribution shows a clear trend: for the reddest LRDs ([0.30.9 m] ), the crossover lies near the Balmer limit, consistent with setton24. For bluer LRDs ([0.30.9 m] ), it shifts redward, often near [O iii]5007 or beyond. Colored markers indicate the individual LRDs shown in Figure 9, illustrating the sequence with color. The bluest example, JADES-30148179, has the longest-wavelength crossover and one of the shallowest optical slopes. First reported by lyu24 as a MIRI-selected AGN, this object, like “Virgil” (iani24), is an example of a blue LRD where the steepening of the optical SED becomes prominent only beyond H , as traced by MIRI.
6 Discussion: Toward a Unified View of the LRD Population
@@
The semi-empirical two-component model discussed in § 5, in which the rest-frame UV continuum is dominated by stellar emission and the optical by a BH⋆, provides a useful framework for interpreting the diversity of LRD SEDs. Combined with the rest-frame color–slope diagram (Figure 3), it serves as a diagnostic tool for exploring how variations in BH and host galaxy properties shape the observed continuum, and potentially tracing evolutionary paths within the LRD population. These ideas are illustrated schematically in Figure 11.
@@
As discussed throughout this work, three key trends emerge in the color–slope diagram when moving from the reddest to the bluest LRDs: (1) The reddest 20% of LRDs ([0.30.9 m] ) exhibit significantly redder UV slopes, while the rest remain nearly constant (); (2) the optical continuum evolves from strong curvature with Balmer breaks and absorption features to smoother, power-law–like shapes; (3) the FWHM of broad Balmer lines become progressively narrower, while narrow metal lines grow stronger and more prominent in the bluest LRDs.
@@
A simple schematic picture (downward arrows in Figure 11) is that the spread in [0.30.9 m] color reflects variations in the relative luminosity of the BH⋆ and host galaxy. The reddest LRDs are dominated by luminous BH⋆ components, while the bluest ones are increasingly galaxy-dominated, not just in the UV but also into the optical (up to 5100 Å), and show stronger narrow lines. As the BH⋆ contribution weakens (due to declining luminosity and/or increased attenuation), the SED inflection point shifts to longer wavelengths, reducing the curvature of the optical continuum and yielding more power-law–like SEDs.
While this picture explains the majority of LRDs with [0.30.9 m] , it raises questions for the reddest ones. What causes their red UV slopes? In our default model, the galaxy dominates the UV (Scenario 1: BH⋆ is UV-opaque), so red UV slopes require higher stellar attenuation and masses (, 7–8.5), unlike the uniformly low values seen in most LRDs. While this could help reduce the extreme ratios in some LRDs, it would imply that the most luminous BH⋆ have predominantly massive and dustier host, at odds with BH⋆ models proposing dense envelopes with little dust attenuation. It also challenges the idea of the evolutionary picture, where the most luminous, fast-growing BH⋆ sources should be hosted by the smallest, youngest galaxies.
@@
An alternative (Scenario 2: BH⋆ is UV-transparent) is that the BH⋆ component emits a weak, heavily reddened UV continuum. Although the dense gas envelope is generally expected to suppress most emission blueward of the Balmer break, certain combinations of gas density and geometry may allow a small amount of reddened UV light to escape (e.g., inayoshi24; degraaff25). In this case, red UV slopes could naturally arise from the BH⋆ rather than requiring dust-reddened hosts. In the bluer LRDs, such BH⋆ UV continuum would be swamped by the galaxy UV emission and unobservable. This scenario allows LRDs to evolve blueward in [0.30.9 m] either via: (1) increasing L/L, or (2) reduced gas density or covering fraction, enabling more UV leakage from the BH⋆.
@@
These two effects could explain UV slope variations among LRDs with similar [0.30.9 m] colors. For example, A2744-45924 and A2744-QSO1 both have [0.30.9 m] 3.5, but very different UV slopes ( and 0; A and C in Figure 9). A2744-45924 may reflect a more transparent BH⋆ with some UV leakage, while A2744-QSO1 could be a “The Cliff”-like system dominated in the UV by a bright, dust-free stellar host (as proposed by naidu25). Other LRDs with similar colors but intermediate slopes (e.g., “Irony” or the Balmer absorbers from kocevski24 and wang24b) likely reflect varying combinations of BH⋆ UV opacity and BH-to-galaxy luminosity ratios. This interpretation naturally produces a sequence where galaxy stellar mass grows from red to blue in the color–slope plane, consistent with a co-evolution scenario in which galaxies become more massive and dust-free as BH⋆ components fade, exiting perhaps a high accretion rate phase or dissolving the dense gas cocoon.

### v2

@@ IV.5 Color Dependence of the Emission Line FWHMs
The colormap suggests a positive correlation, with the reddest LRDs showing the highest FWHMs. A color vs. FWHM linear fit yields a statistically significant correlation, albeit with substantial scatter (Pearson , ; Spearman , ). The median FWHM, binned by rest-frame color as in Figure 5, declines from 2327 760 km s-1 for [0.30.9 m] 3, to 2067 877 km s-1 for [0.30.9 m] , and 1677 955 km s-1 for [0.30.9 m] 2. This trend agrees with the results of greene23, who found broader lines in redder LRDs (e.g., F277WF444W 1.6) in the UNCOVER field.
Despite this general trend, the scatter is large, and there are notable exceptions. Some blue LRDs exhibit very broad lines, for example, GOODS-N-1001830 from juodzbalis24a and UNCOVER-38018 from greene23 both show FWHM 4000 km s-1. Conversely, some of the reddest LRDs, such as RUBIES-UDS-16053 (taylor25) and “The Cliff” (degraaff25), exhibit narrower lines with FWHM 1500 km s-1.
IV.6 LRDs with narrow absorption in the broad emission lines
@@
IV.7 Color Dependence of Emission-line Ratios
@@
The physical origin of the large Balmer decrement in LRDs remains poorly understood. Early interpretations invoking heavy obscuration (e.g., kocevski23, greene23) are disfavored in light of the BH⋆ models, which predict more complex radiative transfer effects. In this context, the elevated broad H /H ratios may be boosted by collisional excitation and resonant scattering in the dense, low-metallicity (and low-extinction) gas cocoon surrounding the accreting black hole (e.g., chang25; torralba25b; yan25). Supporting this scenario, the LRDs with the most extreme H /H ratios are also those that exhibit narrow self-absorption features superimposed on their broad Balmer lines.
@@
To further test whether narrow emission lines arise from the host galaxy while the optical continuum is AGN-dominated, we examine the correlation between equivalent width (EW) and [0.30.9 m] color for the most prominent narrow lines. The right panel of Figure 8 shows the EW of [O iii]5007 and the of blended H +[O iii]4364 the latter serving as a proxy for the complex of lines near the Balmer-break ([O ii], [Ne iii], H ) commonly seen in high-EW galaxies. Among these, H +[O iii]4364 is typically the strongest and most reliably detected, with 45% of LRDs showing S/N 5 in this feature, compared to 20%, 18%, and 1% for [O iii]5007, [Ne iii], and [O ii], respectively. Both EWs show a clear anti-correlation with [0.30.9 m] color: bluer LRDs exhibit stronger narrow-line emission, while redder LRDs show much weaker or undetected lines. This supports the picture that narrow-line features become more prominent toward bluer LRDs, where the host galaxy likely contributes more to the spectrum.
We caution that H +[O iii]4364 may not be purely narrow, especially in the reddest LRDs where broad H emission can dominate. Still, high-resolution NIRSpec spectra suggest that the narrow H component becomes more prominent in bluer LRDs (e.g., akins24b; juodzbalis24b; rinaldi25), while the broad component dominates over the narrow in redder systems (e.g., deugenio25; jones25; torralba25b) similarly to H .
The UV emission lines in LRDs are generally faint, with few detections of classical AGN tracers such as C iii], C iv, or Mg ii above S/N 5. The exception is Ly, detected in 20% of the sample. While Ly equivalent widths show no clear trend with color, detections are concentrated in bluer LRDs, with a median [0.30.9 m] 2.0. Lowering the detection threshold to S/N 2 boosts the C iii] detection rate to 22%, consistent with it being the brightest UV line after Ly; C iv and Mg ii remain below 5%. Like Ly, C iii] detections span a wide color range but are slightly redder on average, with [0.30.9 m] 2.4, consistent with previous reports of C iii] in some of the reddest LRDs (e.g., labbe24b; deugenio25; torralba25b).
V A Semi-Empirical Model for LRD SEDs: From “The Cliff” to “Virgil”
V.1 A Two-Component model: galaxy plus BH⋆
To address these constraints, we adopt a semi-empirical model combining a galaxy and a BH⋆ component. Instead of a complex parametric approach using photoionization codes like CLOUDY (cloudy), we use a simplified method where the BH⋆ component is modeled using the observed continuum of the extreme LRD “The Cliff” (degraaff25). A similar strategy was recently adopted by degraaff25b and umeda25, who approximate the BH⋆ SED with a blackbody. Here, we construct the BH⋆ template by fitting a low-order polynomial to the emission-line–free regions of the “The Cliff” spectrum, capturing its shape from the Balmer break to 1 m rest-frame. We assume no UV emission from the BH⋆ component (see § VI for further discussion). Dust attenuation is applied using an SMC law (pei92), with as the only free parameter modifying the BH⋆ shape. Even modest attenuation can transform the intrinsic curvature of the BH⋆ template into a redder optical continuum that more closely resembles a power law. Varying both the reddening and relative luminosity of the BH⋆ component naturally shifts the SED inflection point to longer wavelengths. Our model does not include an additional IR component to account for potential dust emission from the BH⋆, as in barro25, since the current spectra do not probe the rest-frame NIR (–3 m). Incorporating MIRI data would be essential to assess this contribution and test for warm dust associated with the BH⋆ component.
V.2 Interpreting LRD Diversity Through the Two-Component Model
@@
Our qualitative expectations are confirmed by the distribution of best-fit parameters as a function of rest-frame color across the LRD population. Dividing the sample into three UV-to-optical color bins, [0.30.9 m] , [0.30.9 m] , and [0.30.9 m] , we find that the median stellar attenuation decreases steadily from in the reddest bin to and in the intermediate and bluest bins, respectively. Similarly, the stellar mass declines from to and . In contrast, the attenuation of the BH⋆ component increases across the same bins, from to and . Meanwhile, the BH⋆ luminosity declines from to and . These trends suggest that redder LRDs are powered by more luminous, less obscured BH⋆ components embedded in more attenuated stellar hosts, while bluer LRDs tend to host fainter, more obscured BH⋆ components and nearly dust-free galaxies.
Figure 10 illustrates these trends for the full parameter distributions. For the stellar host properties, both and stellar mass remain nearly constant across the bulk of the population ([0.30.9 m] –3), and then increase toward the reddest, strong-curvature LRDs, accompanied by a larger scatter and uncertainties that reflect the difficulty of constraining the galaxy component when the UV continuum is weak and increasingly dominated by the BH⋆ contribution.
The BH⋆ parameters show a more continuous variation with color. Both the BH⋆ luminosity and evolve smoothly across the full [0.30.9 m] range, but with substantially larger scatter than the stellar properties. In particular, spans a wide range (–4) and exhibits large uncertainties at high attenuation, reflecting the increasing difficulty of constraining the BH⋆ component as it becomes fainter relative to the galaxy in the composite SED. We also highlight LRDs with power-law SEDs (red circles), which preferentially show higher median , consistent with the attenuation required to reproduce their steep optical continua.
@@
Overall, the two-component model successfully fits LRDs across the full range of [0.30.9 m] colors and SED types, suggesting that the observed diversity is primarily driven by variations in the BH⋆-to-galaxy luminosity ratio (L/Lgal), together with increasing obscuration of the BH⋆ component. As L/Lgal declines toward bluer colors, the crossover between the stellar (blue) and BH⋆ (red) continua shifts to longer wavelengths, from 0.4 m to 0.7 m, delaying the onset of the steep optical rise.
@@
To quantify this behavior, Figure 12 shows the wavelength at which the luminosities of the two components are equal, L/L, denoted as . While this crossover does not correspond exactly to the inflection point of the composite SED, the steep slope of the BH⋆ continuum ensures that the turnover typically occurs slightly blueward of . The top y-axis also reports the luminosity ratio L/Lgal at 5100 Å, which provides a measure of the relative normalization of the two components and closely tracks . We find a clear correlation with color: for the reddest LRDs ([0.30.9 m] ), lies near the Balmer limit, consistent with setton24. As [0.30.9 m] decreases and L/Lgal(5100) declines, shifts progressively redward, often approaching [O iii]5007 or longer wavelengths. Red and blue symbols mark the LRDs shown in Figure 11, illustrating the continuous decline with color. For the bluer LRDs (letters E through H), L/Lgal(5100) decreases from 3 to 0.25, indicating that the BH⋆ becomes subdominant at 5100 Å. The bluest example, JADES-30148179, exhibits the longest-wavelength crossover and one of the shallowest optical slopes. First reported by lyu24 as a MIRI-selected AGN, this source, like “Virgil” (iani24), exemplifies blue LRDs in which the steepening of the optical continuum becomes apparent only redward of H , as traced by MIRI photometry.
VI Discussion: Toward a Unified View of the LRD Population
@@
The semi-empirical two-component model discussed in § V, in which the rest-frame UV continuum is dominated by stellar emission and the optical by a BH⋆, provides a useful framework for interpreting the diversity in color and SED shapes of LRD SEDs. Combined with the rest-frame color–slope diagram (Figure 4), it serves as a diagnostic tool for exploring how variations in BH and host galaxy properties shape the observed continuum, and potentially tracing evolutionary paths within the LRD population. These ideas are illustrated schematically in Figure 13.
@@
As discussed throughout this work, three key trends emerge in the color–slope diagram when moving from the reddest to the bluest LRDs: (1) The reddest 20% of LRDs ([0.30.9 m] ) exhibit significantly redder UV slopes, while the majority of the population shows uniformly blue UV slopes with ; (2) the optical continuum shape evolves from strong curvature with Balmer breaks and absorption features to smoother, power-law–like shapes; (3) the FWHM of broad Balmer lines become progressively narrower, while narrow metal lines grow stronger and more prominent in the bluest LRDs.
@@
A simple schematic picture (downward arrows in Figure 13) is that the spread in [0.30.9 m] color reflects variations in the relative luminosity of the BH⋆-to-host (). The reddest LRDs are dominated by luminous BH⋆ components, while the bluest ones are increasingly galaxy-dominated, not just in the UV but also into the optical (up to 5100 Å), and show stronger narrow lines. As declines (due to declining BH⋆ luminosity and/or increased attenuation), the SED inflection point shifts to longer wavelengths, reducing the curvature of the optical continuum and yielding more power-law–like SEDs.
VI.1 What is the Origin of the Red UV Continuum?
While this picture explains the majority of LRDs with [0.30.9 m] , it raises questions for the reddest ones. What causes their red UV slopes? In our default two-component model, the galaxy dominates the UV (Scenario 1: BH⋆ is UV-opaque), so red UV slopes require higher stellar attenuation and masses (–0.5, –9), in contrast to the uniformly lower values observed in most LRDs (, –8). At face value, the higher stellar masses inferred for the reddest LRDs would be consistent with their larger inferred black hole masses, suggested by both higher luminosities and broader Balmer lines, and could help alleviate the tension with the large –0.1 ratios in LRDs relative to local relations (e.g., kocevski23; kocevski24; maiolino23; juodzbalis25; jones25).
However, the preference for dustier hosts is unexpected in the context of dense, dust-poor gas envelopes favored in BH⋆ models for the reddest LRDs. Moreover, the unusual SED of LRDs and the presence of broad wings in their Balmer lines call into question standard bolometric corrections (greene25) and virial linewidth calibrations (rusakov25; kokorev25), potentially lowering inferred MBH by up to an order of magnitude depending on the accretion rates (deugenio25c; torralba25b). Interpreted strictly within Scenario 1, the observed trends would imply an evolutionary sequence from bluer to redder LRDs, with both host galaxies and black holes growing in mass over time.
@@ VI.1 What is the Origin of the Red UV Continuum?
An alternative interpretation (Scenario 2: BH⋆ is UV-transparent) is that the BH⋆ component shows a faint, heavily reddened UV continuum. Although dense gas envelopes are generally expected to suppress emission blueward of the Balmer break, certain combinations of gas density and geometry may allow a small fraction of reddened UV light to escape (e.g., inayoshi24; degraaff25; naidu25). In this case, red UV slopes could arise naturally from the BH⋆ itself, without invoking dusty host galaxies. In bluer LRDs, any such BH⋆ UV emission would be hidden by the brighter galaxy UV continuum.
@@
Within this scenario, shifts toward bluer [0.30.9 m] colors could occur through two channels: (1) decreasing as the host becomes more luminous, or (2) declining gas density or covering fraction, increasing the UV continuum from the BH⋆. Together, these effects can explain the wide range of UV slopes observed at similar [0.30.9 m] colors. For example, A2744-45924 and A2744-QSO1 both have [0.30.9 m] but very different UV slopes ( and ; A and C in Figure 9). A2744-45924 may represent a more transparent BH⋆ with AGN-dominated UV emission (labbe24b), whereas A2744-QSO1 resembles a “Cliff”-like system dominated in the UV by a bright, dust-free stellar host (naidu25). Other LRDs with similar colors but intermediate slopes (e.g., “Irony” or the Balmer absorbers in kocevski24 and wang24b) likely reflect varying combinations of BH⋆ UV opacity and BH-to-galaxy luminosity ratios.
In this second scenario, the reddest LRDs do not require highly attenuated hosts and may instead have lower stellar masses, as indicated by their sub-dominant UV continua. This would imply an evolutionary sequence opposite to that in Scenario 1, with stellar mass increasing from red to blue LRDs. Under this interpretation, the reddest LRDs could correspond to an early phase of rapid, possibly super-Eddington black hole growth in relatively small galaxies. As LRDs evolve toward bluer [0.30.9 m] colors, their host galaxies would grow more massive while the BH⋆ component fades, either by exiting a high accretion phase or by dispersing the dense gas cocoon. A potential challenge to this picture is that black hole masses would need to remain similar or increase toward bluer LRDs, which appears at odds with their lower luminosities and narrower line widths. However, given the uncertainties in interpreting Balmer line profiles in LRDs, a systematic analysis of line shapes across the full population will be essential to robustly test this evolutionary sequence.
VI.2 Is there any Dust in LRDs?
In the two-component model, the higher incidence of power-law–like optical SEDs among bluer LRDs is attributed to increased attenuation of the BH⋆ component, which is the primary parameter capable of modifying the continuum shape. In more physically motivated BH⋆ models, similar reddening could arise from higher gas densities or opacities, reducing the need to invoke dust explicitly. Distinguishing between these possibilities requires assessing whether independent evidence for dust is present.